### v1

@@ 5.1 Neutral Current
Figure 8: Differential distributions of 1-jettiness $\tau$ in a fixed slice of $Q^{2}$ and different bins of $x$ in NC DIS at the EIC. The bottom ratios in each row indicate, from top to bottom, the ratio of each prediction to the MEPS@NLO one, the uncertainty from scale variations and the uncertainty from non-perturbative model and tuning variations. The error bars in the lowest panel indicate the statistical uncertainty of the difference between cluster and string model.
At last, in Fig. 9 we report the 1-jettiness event shape after applying the soft-drop to C ENTAURO jets. We consider three values of the grooming cut from $z_{\text{cut}}\in(0.05,0.1,0.2)$, while remaining inclusive in $Q^{2}$ and $x$. As expected from the removal of soft and low scale radiation, the size of non-perturbative uncertainties is significantly reduced compared to the ungroomed 1-jettiness and does not exhibit strong dependence on $z_{\text{cut}}$. With respect to the MEPS@NLO, the MEPS@LO configuration behaves consistently up to $\tau_{\text{Gr}}<0.5$. Conversely, the MC@NLO and LO+PS ones are suppressed by a factor ${}10\%$. These deviations are still small compared to scale variations, which are reduced switching from MEPS@LO to MEPS@NLO up to $\tau_{\text{Gr}}<0.8$. Moreover, MEPS@NLO scale uncertainties are generally larger than MC@NLO ones, reinforcing the idea in the low-$Q^{2}$ regime LO contributions from higher-multiplicity events provide a large correction.
@@
Figure 9: Groomed 1-jettiness in NC DIS at the EIC. From left to right the soft drop cut $z_{\text{cut}}$ is $0.05$, $0.1$ and $0.2$. The bottom ratios in each row indicate, from top to bottom, the ratio of each prediction to the MEPS@NLO one, the uncertainty from scale variations and the uncertainty from non-perturbative model and tuning variations. The error bars in the lowest panel indicate the statistical uncertainty of the difference between cluster and string model.
@@ 5.2 Charged Current
We again start our discussion with the global observables, $Q^{2}$, $x$, $y$ and missing $p_{T}$ in Figs. 10 and 11. While the inelasticity does not exhibit large differences between the different calculations, we observe rather large corrections for the virtuality and the Bjorken-$x$ distributions. The LO and MC@NLO predictions agree well with each other within the uncertainties, however the merged setups receive large corrections towards decreasing $Q^{2}$, $x$ and missing $p_{T}$. This trend starts at values of $10^{3}\,\text{GeV}^{2}$ and $10^{-1}$, respectively, and reach up to almost a factor of 2 for the lowest $Q^{2}$ and $x$ values. With respect to each other the MEPS@LO and MEPS@NLO agree well, the uncertainties are significantly reduced in the latter, as already observed before. Naturally, the MEPS@NLO shows smaller uncertainties than the MEPS@LO and comparable ones to the MC@NLO at high values for the three observables. As multi-jet events dominate at low $Q^{2}$ and these events are only computed at LO, the MEPS@NLO shows larger uncertainties than MC@NLO there. Non-perturbative uncertainties were negligible for these observable as they are measured by means of the lepton kinematics and are hence not affected by hadronization.
Figure 10: Differential distributions for $Q^{2}$ and $x$ in Charged Current DIS at the EIC, comparing LO+PS, MC@NLO, MEPS@LO and MEPS@NLO predictions. The main panels show results with scale variation uncertainties for all but the LO+PS calculation. The first ratio panel displays the relative deviation from MEPS@NLO to highlight merging effects. The second ratio shows the relative scale uncertainties, while the third presents the breakdown of hadronization uncertainties as detailed in Sec. 3.2.
Figure 11: Differential distributions for $y$ and $p_{\text{T}}^{\mathrm{missing}}$ in Charged Current DIS at the EIC, comparing LO+PS, MC@NLO, MEPS@LO, and MEPS@NLO. Main panels show scale variation uncertainties (excluding LO+PS). Ratio panels display merging effects and scale uncertainties.
[…]
@@
Figure 12: Jet multiplicity $N_{\mathrm{jet}}$ (top row) and leading jet transverse momentum $p_{T}^{\mathrm{jet}}$ (bottom row) distributions for anti-$k_{t}$ (left), $k_{t}$ (middle) and C ENTAURO (right) jet algorithms in CC DIS at the EIC. The bottom ratios in each row indicate, from top to bottom, the ratio of each prediction to the MEPS@NLO one, the uncertainty from scale variations and the uncertainty from non-perturbative model and tuning variations. The error bars in the lowest panel indicate the statistical uncertainty of the difference between cluster and string model.
Analogously to before, we show the 1-jettiness in bins of $Q^{2}$ and $x$ in Fig. 13 without ratio plots for clarity and for the range $10031.6<Q^{2}/\,\text{GeV}^{2}<$ 100 in three $x$ ranges including ratio plots in Fig. 14. Note that we pick a different $Q^{2}$ slice with respect to the NC case, as the total cross-section in the lower slice is smaller. Plots for the remaining $x$-$Q^{2}$ ranges can again be found in App. A.2. Overall, we see similar changes in the shape when going from small to large virtualities. As the calculation of the 1-jettiness operates inclusively, i.e. without a jet cut, we can observe large corrections on the cross-section for low $Q^{2}$ by merging additional legs. With regard to the hadronization, we also see the same pattern as before, i.e. only the peak is affected and the phase space region of small values of $x$ and $Q^{2}$ has larger uncertainties. It is only in these regions that the perturbative and the non-perturbative uncertainties are of comparable size. While 1-jettiness is only an example, this suggests that global analyses of charged-current events at the EIC need higher order corrections for the small virtualities. Depending on the observable, a thorough understanding of non-perturbative effects will also be necessary.
Figure 13: Differential distributions of 1-jettiness $\tau$ in different bins of $Q^{2}$ and $x$ in CC DIS at the EIC. The total MEPS@NLO cross-section for each $x$-$Q^{2}$ bin is reported on top of the corresponding plot.
Figure 14: Differential distributions of 1-jettiness $\tau$ in a fixed slice of $Q^{2}$ and different bins of $x$ in Charged Current DIS at the EIC. The bottom ratios in each row indicate, from top to bottom, the ratio of each prediction to the MEPS@NLO one, the uncertainty from scale variations and the uncertainty from non-perturbative model and tuning variations. The error bars in the lowest panel indicate the statistical uncertainty of the difference between cluster and string model.
@@ 6 Conclusion
In this work we presented the first predictions at MEPS@NLO precision for both neutral and charged current DIS at the EIC. Producing full hadron-level predictions, we contrasted perturbative and non-perturbative uncertainties and identified regions where higher-order corrections will be necessary. To this end, we used the multi-jet merging capabilities of the S HERPA event generator, which lead to non-negligible corrections in the cross-sections compared to LO. In the analysis we focused on global DIS observables, jet multiplicities and transverse momenta in the Breit frame and 1-jettiness distributions. While estimating the perturbative higher order uncertainties with 7-point scale variations we examined the non-perturbative uncertainties associated with the fragmentation modeling by means of replica tunes of the hadronization and the uncertainties associated with the choice of non-perturbative model.
@@
We found that the merging of additional legs in the matrix-element in the MEPS@LO and MEPS@NLO schemes, respectively, captures large corrections of up to factor 2 in the low $Q^{2}$ and $x$ phase space. The virtual corrections lead to an observable effect on the cross-section, especially at high $Q^{2}$ and $x$. Comparing the $k_{t}$, anti-$k_{t}$ and C ENTAURO jet algorithms, we observed very large $K$-factors for the two former, but only a small $K$-factor for the latter. We assume this to be a consequence of the C ENTAURO algorithm being better at clustering back to a Born kinematics, whereas only higher order corrections contribute for $k_{t}$-type algorithms in the Breit frame.
@@
We then studied 1-jettiness as an example for a global event-shape observable, enabling among others the extraction of the strong coupling in measurements at the EIC. We analyzed it in slices of $Q^{2}$ and $x$ and observed a change in event shapes from single-jet topologies at high virtualities towards more uniform shapes at small virtualities. Non-perturbative uncertainties can reach comparable size as the scale variations in limited region of the phase space and especially towards smaller virtualities. Those can to an extent be mitigated by modern grooming techniques. However, especially with the availability of NNLO FO calculations, and the advent of general NNLO matching techniques in the foreseeable future, this implies a strong necessity of understanding better non-perturbative and higher twist corrections due to the lower center-of-mass energy at the EIC than at HERA, in order to fully exploit the potential of the data. Tuning cluster and string model in comparable setups, using simultaneously LEP and DIS data, should be a pragmatic first step in this direction. Ideally, this exercise would also produce tuning uncertainties for both model tunes.
@@
Our study showcases the state-of-the-art of precision hadron-level predictions for the EIC and examines the leading uncertainties. At virtualities of $Q^{2}\approx 1\text{-}10\,\text{GeV}^{2}$ in NC DIS, the photoproduction region comes into play, and the merging of additional legs in the matrix elements are the tool to interpolate towards this region.
@@
While many exclusive measurements are planned during the operation of the EIC, an understanding of the inclusive background is crucial for any analysis. Here, we showed a global analysis of the DIS region, future studies will need to understand better the photoproduction region and the interpolation between the two to fully allow inclusive jet measurements at the EIC. This would allow performing jet studies at the EIC in a fully inclusive way over all virtualities, which has so far been formulated as a goal in the experimental program, but has not seen much attention from the theoretical side. Calculations at even higher order, at NNLO and beyond, and ideally combined with either resummation or matched to the parton shower, should be within reach in the upcoming years and in time for the data-taking at the EIC.
@@ Acknowledgments
We are grateful to Frank Krauss for encouraging us to investigate this topic and for his comments on our manuscript. D.R. is supported by the European Union under the HORIZON program in Marie Skłodowska-Curie project No. 101153541. F.S. is supported by the STFC under grant agreement ST/P006744/1. P.M. is supported by the Swiss National Science Foundation (SNF) under contract 200020-204200.
@@ A.1 Neutral current 1-jettiness
Here we show the remaining 1-jettiness histograms in NC DIS in the $x$ and $Q^{2}$ ranges from Fig. 7 with the same ratio panels as in Fig. 8.
@@
Figure 15: Differential distributions of 1-jettiness $\tau$ in a fixed slice of $Q^{2}$ and different bins of $x$ in NC DIS at the EIC. The bottom ratios in each row indicate, from top to bottom, the ratio of each prediction to the MEPS@NLO one, the uncertainty from scale variations and the uncertainty from non-perturbative model and tuning variations. The error bars in the lowest panel indicate the statistical uncertainty of the difference between cluster and string model.
@@
Figure 16: Differential distributions of 1-jettiness $\tau$ in a fixed slice of $Q^{2}$ and different bins of $x$ in NC DIS at the EIC. The bottom ratios in each row indicate, from top to bottom, the ratio of each prediction to the MEPS@NLO one, the uncertainty from scale variations and the uncertainty from non-perturbative model and tuning variations. The error bars in the lowest panel indicate the statistical uncertainty of the difference between cluster and string model.
@@ A.2 Charged current 1-jettiness
Here we show remaining the 1-jettiness observable in CC DIS in the $x$ and $Q^{2}$ ranges from Fig. 13 with the same ratio panels as in Fig. 14.
@@
Figure 17: Differential distributions of 1-jettiness $\tau$ in a fixed slice of $Q^{2}$ and different bins of $x$ in CC DIS at the EIC. The bottom ratios in each row indicate, from top to bottom, the ratio of each prediction to the MEPS@NLO one, the uncertainty from scale variations and the uncertainty from non-perturbative model and tuning variations. The error bars in the lowest panel indicate the statistical uncertainty of the difference between cluster and string model.
@@
Figure 18: Differential distributions of 1-jettiness $\tau$ in a fixed slice of $Q^{2}$ and different bins of $x$ in CC DIS at the EIC. The bottom ratios in each row indicate, from top to bottom, the ratio of each prediction to the MEPS@NLO one, the uncertainty from scale variations and the uncertainty from non-perturbative model and tuning variations. The error bars in the lowest panel indicate the statistical uncertainty of the difference between cluster and string model.

### v2

@@ 5.1 Neutral current
Figure 8: Differential distributions of 1-jettiness $\tau$ in a fixed slice of $Q^{2}$ and different bins of $x$ in NC DIS at the EIC. The bottom ratios in each row indicate, from top to bottom, the ratio of each prediction to the MEPS@NLO one, the uncertainty from scale variations and the uncertainty from nonperturbative model and tuning variations. The error bars in the lowest panel indicate the statistical uncertainty of the difference between cluster and string model.
At last, in Fig. 9 we report the 1-jettiness event shape after applying the soft drop to C ENTAURO jets. We consider three values of the grooming cut from $z_{\text{cut}}\in(0.05,0.1,0.2)$, while remaining inclusive in $Q^{2}$ and $x$. As expected from the removal of soft and low scale radiation, the size of nonperturbative uncertainties is significantly reduced compared to the ungroomed 1-jettiness and does not exhibit strong dependence on $z_{\text{cut}}$. With respect to the MEPS@NLO, the MEPS@LO configuration behaves consistently up to $\tau_{\text{Gr}}<0.5$. Conversely, the MC@NLO and LO+PS ones are suppressed by a factor $10\%$. These deviations are still small compared to scale variations, which are reduced switching from MEPS@LO to MEPS@NLO up to $\tau_{\text{Gr}}<0.8$. Moreover, MEPS@NLO scale uncertainties are generally larger than MC@NLO ones, reinforcing the idea in the low-$Q^{2}$ regime LO contributions from higher-multiplicity events provide a large correction.
@@
Figure 9: Groomed 1-jettiness in NC DIS at the EIC. From left to right the soft drop cut $z_{\text{cut}}$ is $0.05$, $0.1$ and $0.2$. The bottom ratios in each row indicate, from top to bottom, the ratio of each prediction to the MEPS@NLO one, the uncertainty from scale variations and the uncertainty from nonperturbative model and tuning variations. The error bars in the lowest panel indicate the statistical uncertainty of the difference between cluster and string model.
@@ 5.2 Charged current
We again start our discussion with the global observables, $Q^{2}$, $x$, $y$ and missing $p_{T}$ in Figs. 10 and 11. While the inelasticity does not exhibit large differences between the different calculations, we observe rather large corrections for the virtuality and the Bjorken-$x$ distributions. The LO and MC@NLO predictions agree well with each other within the uncertainties; however, the merged setups receive large corrections toward decreasing $Q^{2}$, $x$ and missing $p_{T}$. This trend starts at values of $10^{3}\,\text{GeV}^{2}$ and $10^{-1}$, respectively, and reach up to almost a factor of 2 for the lowest $Q^{2}$ and $x$ values. With respect to each other the MEPS@LO and MEPS@NLO agree well, and the uncertainties are significantly reduced in the latter, as already observed before. Naturally, the MEPS@NLO shows smaller uncertainties than the MEPS@LO and comparable ones to the MC@NLO at high values for the three observables. As multijet events dominate at low $Q^{2}$ and these events are only computed at LO, the MEPS@NLO shows larger uncertainties than MC@NLO there. Likewise, merging scale variations are considerable in this range, analogous to our discussion for the NC case. nonperturbative uncertainties were negligible for these observable as they are measured by means of the lepton kinematics and are hence not affected by hadronization.
Figure 10: Differential distributions for $Q^{2}$ and $x$ in charged current DIS at the EIC, comparing LO+PS, MC@NLO, MEPS@LO and MEPS@NLO predictions. The main panels show results with scale variation uncertainties for all but the LO+PS calculation. The first ratio panel displays the relative deviation from MEPS@NLO to highlight merging effects. The second ratio shows the relative scale uncertainties, as detailed Sec. 3.2. Full (hatched) band for $\mu_{\text{F}},\mu_{\text{R}}$ ($Q_{\mathrm{cut}}$) respectively.
Figure 11: Differential distributions for $y$ and $p_{\text{T}}^{\mathrm{missing}}$ in charged current DIS at the EIC, comparing LO+PS, MC@NLO, MEPS@LO, and MEPS@NLO. Main panels show scale variation uncertainties (excluding LO+PS). Ratio panels display merging effects (hatched band) and scale uncertainties (full bands).
In Fig. 12, we study the three different jet algorithms in terms of jet multiplicity and leading jet transverse momentum. In the latter, and unlike to the neutral current case, there are no corrections with respect to the LO calculation in the low-$p_{T}$ region for neither of the clustering algorithms. At high $p_{T}$ the simulation at LO accuracy yields a slightly lower cross section by about 50% compared to the MEPS@NLO result while the other calculations agree with each other. Hence, we can assume that the large corrections that can be seen in $Q^{2}$ and $x$ are driven by small-$p_{T}$ jets, which are removed by the $p_{T}$ cut. Seeing that already MC@NLO does not show large differences to the merged setups, we conclude that for jet observables with the used $p_{T}$ cut of $5\,\text{GeV}$, already the real correction of the Born process captures any corrections in this part of the phase space. As a consequence, the scale uncertainties of the MC@NLO and the MEPS@NLO do not differ as much as in the neutral current case before. Merging scale variations are again less important than scale variations. As in the NC case, nonperturbative uncertainties are generally smaller than perturbative ones and affect significantly only the high-multiplicity and low-$p_{T}$ regions. At most, they account for a relative $\sim 5\%$ correction, which matches the one induced by scale variations only at very low $p_{T}$ for all jet algorithms considered. Fragmentation model uncertainty is the dominant contribution, while replica tunes within the cluster model do not generate an observable effect within statistical fluctuations.
@@
Figure 12: Jet multiplicity $N_{\mathrm{jet}}$ (top row) and leading jet transverse momentum $p_{T}^{\mathrm{jet}}$ (bottom row) distributions for anti-$k_{t}$ (left), $k_{t}$ (middle) and C ENTAURO (right) jet algorithms in CC DIS at the EIC. The bottom ratios in each row indicate, from top to bottom, the ratio of each prediction to the MEPS@NLO one, the uncertainty from scale variations (full band for $\mu_{\text{F}}\,\mathrm{and}\,\mu_{\text{R}}$, hatched for $Q_{\mathrm{cut}}$) and the uncertainty from nonperturbative model and tuning variations. The error bars in the lowest panel indicate the statistical uncertainty of the difference between cluster and string model.
Analogously to before, we show the 1-jettiness in bins of $Q^{2}$ and $x$ in Fig. 13 without ratio plots for clarity and for the range $31.6<Q^{2}/\,\text{GeV}^{2}<100$ in three $x$ ranges including ratio plots in Fig. 14. Note that we pick a different $Q^{2}$ slice with respect to the NC case, as the total cross section in the lower slice is smaller. Plots for the remaining $x$-$Q^{2}$ ranges can again be found in Appendix A.2. Overall, we see similar changes in the shape when going from small to large virtualities. As the calculation of the 1-jettiness operates inclusively, i.e. , without a jet cut, we can observe large corrections on the cross section for low $Q^{2}$ by merging additional legs. With regard to the hadronization, we also see the same pattern as before; i.e. , only the peak is affected and the phase space region of small values of $x$ and $Q^{2}$ has larger uncertainties. It is only in these regions that the perturbative and the nonperturbative uncertainties are of comparable size. While 1-jettiness is only an example, this suggests that global analyses of charged-current events at the EIC need higher-order corrections for the small virtualities. Depending on the observable, a thorough understanding of nonperturbative effects will also be necessary.
Figure 13: Differential distributions of 1-jettiness $\tau$ in different bins of $Q^{2}$ and $x$ in CC DIS at the EIC. The total MEPS@NLO cross section for each $x$-$Q^{2}$ bin is reported on top of the corresponding plot.
Figure 14: Differential distributions of 1-jettiness $\tau$ in a fixed slice of $Q^{2}$ and different bins of $x$ in charged current DIS at the EIC. The bottom ratios in each row indicate, from top to bottom, the ratio of each prediction to the MEPS@NLO one, the uncertainty from scale variations and the uncertainty from nonperturbative model and tuning variations. The error bars in the lowest panel indicate the statistical uncertainty of the difference between cluster and string model.
@@ 6 Conclusion
In this work we presented the first predictions at MEPS@NLO precision for both neutral and charged current DIS at the EIC. Producing full hadron-level predictions, we contrasted perturbative and nonperturbative uncertainties and identified regions where higher-order corrections will be necessary. To this end, we used the multijet merging capabilities of the S HERPA event generator, which lead to non-negligible corrections in the cross sections compared to LO. In the analysis we focused on global DIS observables, jet multiplicities, and transverse momenta in the Breit frame and 1-jettiness distributions. While estimating the perturbative higher-order uncertainties with seven-point scale variations, we examined the nonperturbative uncertainties associated with the fragmentation modeling by means of replica tunes of the hadronization and the uncertainties associated with the choice of nonperturbative model.
@@
We found that the merging of additional legs in the matrix element in the MEPS@LO and MEPS@NLO schemes, respectively, captures large corrections of up to factor 2 in the low $Q^{2}$ and $x$ phase space. The virtual corrections lead to an observable effect on the cross section, especially at high $Q^{2}$ and $x$. Comparing the $k_{t}$, anti-$k_{t}$ and C ENTAURO jet algorithms, we observed very large $K$-factors for the two former, but only a small $K$-factor for the latter. We assume this to be a consequence of the C ENTAURO algorithm being better at clustering back to a Born kinematics, whereas only higher-order corrections contribute for $k_{t}$-type algorithms in the Breit frame.
@@
We then studied 1-jettiness as an example for a global event-shape observable, enabling among others the extraction of the strong coupling in measurements at the EIC. We analyzed it in slices of $Q^{2}$ and $x$ and observed a change in event shapes from single-jet topologies at high virtualities toward more uniform shapes at small virtualities ones. nonperturbative uncertainties can reach comparable size as the scale variations in limited region of the phase space and especially toward smaller virtualities. Those can to an extent be mitigated by modern grooming techniques. However, especially with the availability of NNLO FO calculations, and the advent of general NNLO matching techniques in the foreseeable future, this implies a strong necessity of understanding better nonperturbative and higher-twist corrections due to the lower center-of-mass energy at the EIC than at HERA, in order to fully exploit the potential of the data. Tuning cluster and string model in comparable setups, using simultaneously LEP and DIS data, should be a pragmatic first step in this direction. Ideally, this exercise would also produce tuning uncertainties for both model tunes.
@@
Our study showcases the state-of-the-art of precision hadron-level predictions for the EIC and examines the leading uncertainties. At virtualities of $Q^{2}\approx 1\text{-}10\,\text{GeV}^{2}$ in NC DIS, the photoproduction region comes into play, and the merging of additional legs in the matrix elements are the tool to interpolate toward this region.
@@
While many exclusive measurements are planned during the operation of the EIC, an understanding of the inclusive background is crucial for any analysis. Here, we showed a global analysis of the DIS region, future studies will need to understand better the photoproduction region and the interpolation between the two to fully allow inclusive jet measurements at the EIC. This would allow performing jet studies at the EIC in a fully inclusive way over all virtualities, which has so far been formulated as a goal in the experimental program, but has not seen much attention from the theoretical side. Calculations at even higher order, at NNLO and beyond, and ideally combined with either resummation or matched to the parton shower, should be within reach in the upcoming years and in time for the data taking at the EIC.
@@ Acknowledgments
We are grateful to Frank Krauss for encouraging us to investigate this topic and for his comments on our manuscript. D.R. is supported by the European Union under the HORIZON program in Marie Skłodowska-Curie Project No. 101153541. F.S. is supported by the STFC under Grant No. ST/P006744/1. P.M. is supported by the Swiss National Science Foundation (SNF) under Contract No. 200020-204200.
@@ A.1 Neutral current 1-jettiness
Here, we show the remaining 1-jettiness histograms in NC DIS in the $x$ and $Q^{2}$ ranges from Fig. 7 with the same ratio panels as in Fig. 8. These plots are shown in Figs. 15 and 16.
@@
Figure 15: Differential distributions of 1-jettiness $\tau$ in a fixed slice of $Q^{2}$ and different bins of $x$ in NC DIS at the EIC. The bottom ratios in each row indicate, from top to bottom, the ratio of each prediction to the MEPS@NLO one, the uncertainty from scale variations and the uncertainty from nonperturbative model and tuning variations. The error bars in the lowest panel indicate the statistical uncertainty of the difference between cluster and string model.
@@
Figure 16: Differential distributions of 1-jettiness $\tau$ in a fixed slice of $Q^{2}$ and different bins of $x$ in NC DIS at the EIC. The bottom ratios in each row indicate, from top to bottom, the ratio of each prediction to the MEPS@NLO one, the uncertainty from scale variations and the uncertainty from nonperturbative model and tuning variations. The error bars in the lowest panel indicate the statistical uncertainty of the difference between cluster and string model.
@@ A.2 Charged current 1-jettiness
Here, we show remaining the 1-jettiness observable in CC DIS in the $x$ and $Q^{2}$ ranges from Fig. 13 with the same ratio panels as in Fig. 14. These plots are shown in Figs. 17 and 18.
@@
Figure 17: Differential distributions of 1-jettiness $\tau$ in a fixed slice of $Q^{2}$ and different bins of $x$ in CC DIS at the EIC. The bottom ratios in each row indicate, from top to bottom, the ratio of each prediction to the MEPS@NLO one, the uncertainty from scale variations and the uncertainty from nonperturbative model and tuning variations. The error bars in the lowest panel indicate the statistical uncertainty of the difference between cluster and string model.
@@
Figure 18: Differential distributions of 1-jettiness $\tau$ in a fixed slice of $Q^{2}$ and different bins of $x$ in CC DIS at the EIC. The bottom ratios in each row indicate, from top to bottom, the ratio of each prediction to the MEPS@NLO one, the uncertainty from scale variations and the uncertainty from nonperturbative model and tuning variations. The error bars in the lowest panel indicate the statistical uncertainty of the difference between cluster and string model.